Convert the following binary numbers to decimal and show your calculations: 10101100011 and 10011111
Convert the following decimal numbers to binary and show your calculations: 7625 and 872

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

Convert the following binary numbers to decimal and show your calculations: 10101100011
= 563 in Hex
= 5*256 + 6*16 + 3 = 1280 + 96 + 3
= 1379 decimal
-------------------------
and 10011111
= 9F in Hex
= 9*16 + 15
= 159 decimal
===========================
Convert the following decimal numbers to binary and show your calculations:
7625 - 4096 = 3529 --> 1xxx
INT(3529/256) = 13 --> 1Dxx
(Remainder * 16) = 12 --> 1DCx
Remainder = 9 --> 1DC9 in Hex
--> 1 1101 1100 1001 in Binary
=================================
and 872
872 - 256*3 = 104 --> 3xx
104 - 16*6 = 8 --> 36x
--> 368 Hex
= 0011 0110 1000 in Binary
=================================
The best news is Octal (base 8) seems to have blown over.

Answer by Edwin McCravy(20054)

I think that may be too advanced for you.  Here's the more
elementary way to do the conversions.

Convert the following binary numbers to decimal and show your
calculations: 10101100011 and 10011111
List the digits in a vertical column, each followed by " + 2×"    

1 + 2×
0 + 2×
1 + 2×
0 + 2×
1 + 2×
1 + 2×
0 + 2×
0 + 2×
0 + 2×
1 + 2×
1 + 2×

Begin with a 0 after the × on the first row, then an equal sign, 
then do the calculation and put the answer after another equal sign:

1 + 2×0 = 1 + 0 = 1 
0 + 2×
1 + 2×
0 + 2×
1 + 2×
1 + 2×
0 + 2×
0 + 2×
0 + 2×
1 + 2×
1 + 2×

Then put that result, 1, after the × on the 2nd row, 
then do that calculation and put that answer after 
another equal sign:

1 + 2×0 = 1 + 0 = 1 
0 + 2×1 = 0 + 2 = 2 
1 + 2×
0 + 2×
1 + 2×
1 + 2×
0 + 2×
0 + 2×
0 + 2×
1 + 2×
1 + 2×
 
Then put that result, 2, after the × on the 3rd row, 
then do that calculation and put that answer after 
another equal sign:

1 + 2×0 = 1 + 0 = 1 
0 + 2×1 = 0 + 2 = 2 
1 + 2×2 = 1 + 4 = 5
0 + 2×
1 + 2×
1 + 2×
0 + 2×
0 + 2×
0 + 2×
1 + 2×
1 + 2×

Keep doing that all the way down to the bottom:

1 + 2×0 = 1 + 0 = 1 
0 + 2×1 = 0 + 2 = 2 
1 + 2×2 = 1 + 4 = 5
0 + 2×5 = 0 + 10 = 10 
1 + 2×10 = 1 + 20 = 21 
1 + 2×21 = 1 + 42 = 43 
0 + 2×43 = 0 + 86 = 86
0 + 2×86 = 0 + 172 = 172
0 + 2×172 = 0 + 344 = 344
1 + 2×344 = 1 + 688 = 689 
1 + 2×689 = 1 + 1378 = 1379

Answer = 1379

Now you do the other one the same way.

Convert the following decimal numbers to binary and show your
calculations: 7625
Start by dividing 7625 by 2, getting 3812 with remainder 1 
and placing the remainder R=1 out beside the quotient.
Then divide the quotient 3812 by 2, getting 1906 with remainder 0
and placing the quotiont R=0 out beside the quotient.
Do that all the way down until the quotient is 0.

2)7625
2)3812 R=1
2)1906 R=0
 2)953 R=0
 2)476 R=1
 2)238 R=0
 2)119 R=0
  2)59 R=1
  2)29 R=1
  2)14 R=1
   2)7 R=0
   2)3 R=1
   2)1 R=1
   2)0 R=1      

Now take those remainders in reverse order,
that is, from the bottom to the top:

1110111001001

That's the answer.

Now you do the other one the same way. 

Edwin